OpenStudy (anonymous):

How do you solve this problem with a fraction answer. 3=k(sqr 2.5) I get 1.603567451475 and type 1.6 as my answer but get it wrong so what would be the fraction version to see if the online homework just simply wants a fraction?

OpenStudy (anonymous):

the above equation is actually \[3 = k \sqrt{3.5}\]

OpenStudy (wildweasel):

The key thing here would be to recognize that you can express 3.5 as a fraction, specifically, 7/2. Use that instead of 3.5 and solve like you normally would. You should get a fraction answer.

OpenStudy (anonymous):

I get \[\frac{ 3\sqrt{14} }{ 7 }\]

OpenStudy (wildweasel):

To check your answer, just plug in your result to k and see if the left side equals the right side of the equation. If it does (which it should), you've done it right. With online homework questions, its tough to tell sometimes exactly what theyre looking for because there are many equivalent representations of the answer. For example, when I did this problem, the answer I got was \[3\sqrt{\frac{ 2 }{ 7 }}\] which is equivalent to your solution.